OpenStudy (anonymous):

let f(x)=x^2-100.the average rate of change from x=-2 to x=2 equals A.-100 B.100 C.0 D.400

OpenStudy (anonymous):

@SithsAndGiggles can u help me solve his please

OpenStudy (ipwnbunnies):

The average rate of change will be equivalent to the slope of the line that connects the two points at x=2 and x=-2

OpenStudy (ipwnbunnies):

So, use the slope formula. But, you'll have to find the values of the function at x=-2 and x=2

OpenStudy (anonymous):

ok can u help me figure out the answer

OpenStudy (ipwnbunnies):

Sure. Find the value of the function at x=-2. Plug -2 in for x.

OpenStudy (anonymous):

-104

OpenStudy (ipwnbunnies):

Try again. y = (-2)^2 - 100

OpenStudy (anonymous):

i got the same thing can u help

OpenStudy (ipwnbunnies):

(-2)^2 = 4 <- what we're using -2^2 = -4 <- not correct We're squaring the entire x value of -2. (-2)^2 = -2*-2 = 4

OpenStudy (ipwnbunnies):

When x=-2, y = (-2)^2 - 100 = 4 - 100 = -96

OpenStudy (anonymous):

ok

OpenStudy (ipwnbunnies):

Find the y-value at x=2. now.

OpenStudy (anonymous):

2

OpenStudy (ipwnbunnies):

What. ;-; Plug in 2 for x, into the function.

OpenStudy (anonymous):

need help

OpenStudy (ipwnbunnies):

f(x) = x^2 - 100 That's the function you were given. .-. f(2) = 2^2 - 100

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

s its -100

OpenStudy (ipwnbunnies):

.-.

OpenStudy (ipwnbunnies):

f(2) = 2^2 - 100 = 4 - 100 = ?

OpenStudy (ipwnbunnies):

Then use the slope formula. This will give you the slope of the line connecting the two points. That slope is equivalent to the average rate of change. \[slope = \frac{y_{2} - y_{1}}{x_{2} - x_{1}}\]